# Simple portable quantum key distribution for science outreach

### v1

@@ I Introduction
While educational articles exist on QKD implementation in laboratories educational , here we focus on the use of weak coherent pulses which offer a more straightforward and cost-effective setup, making quantum communication protocols accessible to undergraduate students. This work aims to bridge the gap between theory and practice, providing a valuable educational resource for understanding and implementing secure quantum communication.
@@ II.2 One Time Pad and BB84
Bob chooses one measurement basis and performs the measurement. For the same choice of basis, Alice and Bob will share the same bit. For example, if Alice sends bit 0 in the Z basis and Bob performs a measurement on that basis, the probability of Bob measuring bit 0 is given by
@@
This means that Bob cannot recover the correct bit string when measuring on a different basis every round. However, this measurement error from choosing different bases can help us identify the presence of Eve. If Eve intercepts and resends the state she measured from Alice she will have to introduce errors in the communicated message.
As the eavesdropper (usually referred to as Eve) does not know which basis both parties chose, she will guess a measurement basis. If she chooses a different basis from both parties, she will prepare a different state than the original and introduce errors.
@@ II.3 Simplified BB84
For educational purposes focused on understanding secure communication through Quantum Key Distribution (QKD), error correction and privacy amplification are not relevant. These steps are typically used to correct errors and enhance security but are not essential for grasping the fundamental concepts of QKD implementation.
For this protocol, the security analysis and parameter estimation changes, and the details can be found in .QBER
@@ III Experimental implementation
This setup is divided into two parts, two optical breadboards of 60x60 cm (Thorlabs, B6060A), the emitter, and the receiver. These are the breadboards available in the laboratory but smaller breadboards could be used instead. The emitter is capable of sending weak coherent pulsesCoherent in three different polarization states ($|\mathrm{H}\rangle,\ |\mathrm{V}\rangle,\ |\mathrm{D}\rangle$). The receiver in turn is capable of measuring on two different bases, $Z$ and $X$.
@@ III.1 Emitter
The purple pulses are utilized to modulate the VCSEL to switch the laser on and off. The synchronization shown in the figure demonstrates the simultaneous modulation of these two types of pulses: blue for polarization control and purple for laser activation. This synchronization enables independent polarization control for each light pulse. In this case, the pulse period was 10 $\mu s$ and the repetition rate was 50 kHz.
@@
Alternatively, it is also possible to implement a less costly approach, removing the EOM from the setup, to the emitter as seen in Fig. 3. Instead of one laser, three would be used. Only one laser would be turned on at any time and the path taken by the light would create the respective state. To create the $|\mathrm{H}\rangle$ and $\ |\mathrm{V}\rangle$ states, only a PBS would be required. To create the $|\mathrm{D}\rangle$ state, a PBS followed by a HWP can be used. This way, the state choice would be done by choosing which laser to turn on instead of by active modulation of the light. The different paths would then be recombined using a 50/50 beam splitter (BS)(Thorlabs, BS011) and the rest of the emitter would work as in the active approach.
This method brings security risks as the lasers can have manufacturing differences (spatial mode, frequency…) and an adversary might take advantage of those but for this work, it can be a solution.
@@
Given that both the emitter and receiver are situated within the same room, the free-space channel is defined as the distance between the telescopes at each end. While a well-collimated beam and a short free-space channel may render the use of telescopes optional, their inclusion enables easy scalability of the channel length by adjusting the distance between optical tables.
@@ III.2 Receivers
The objective of the receiver is to measure the incoming states on one of two basis. By calculating the Quantum Bit Error Rate (QBER), we can detect the presence of potential eavesdroppers (Eve) and generate secret keys. For educational purposes focusing on the concept of key generation rather than actual key generation, this process can be simplified.
To accomplish this, as seen in Fig. 1, the light is captured by a telescope and directed through a HWP, followed by a PBS. The angle in the HWP will define if the measure is on the Z basis or the X basis. Each output of the PBS is then coupled to a fiber and a single-photon detector (SPD) (Excelitas, SPCM-AQRH-10). The arrival time of the photons is recorded using a time tagger (Swabian Instruments, Time Tagger 20), which offers precision beyond the requirement; a more affordable option such as a microcontroller with a 100 MHz timer frequency (10 ns accuracy) can be used. Additionally, the trigger pulse modulating the laser should be connected to the time tagger for post-processing of the measured data.
@@ III.3 Parameter definition
In a secure implementation of the protocol, the average number of photons emitted from the emitter should be low. However, for educational purposes, we may increase this value to simplify detection and result analysis. This allows for a higher loss in the receiver while still ensuring an adequate average number of photons to generate sufficient events for straightforward post-processing.
@@
From Eq.(4), the necessary filter loss in dB can be found.
@@
To achieve this with only two detectors, measurements on the X basis are deliberately delayed. Despite the lack of refractive index information for the fiber used, typical values generally fall within the range of [1.45, 1.48] $\mu$s per meter. Utilizing an available 1590-meter optical fiber, although not ideal, yields an expected delay of approximately [7.685, 7.844] $\mu$s.
To verify this delay and assess time-multiplexed detection, a low-photon-count pulse with diagonal polarization is emitted from the emitter. For experimental validation, the BS in the setup (Fig 4) is replaced with a PBS to ensure minimal overlap between delayed and not delayed beams. The horizontal component of the polarized beam traverses a second PBS, reaching only detector 1 after passing through a filter to simulate losses in the alternate path. Conversely, the vertical component is directed into the optical fiber for temporal delay. Subsequently, a HWP rotates the polarization at the fiber output to horizontal, guaranteeing that photons exclusively reach detector 2 via a second PBS.
Consequently, clicks registered in detector 1 originate from the not delayed signal, while those in detector 2 result from the delayed signal.
@@
The overlap observed in the distributions can be attributed to selecting a delay shorter than the pulse width. However, the overlap is minimal, allowing for clear differentiation between not delayed and delayed signals and facilitating the measurement of states in different bases. Additionally, a broadening of the delayed beam is evident, likely caused by dispersion in the fiber.
@@ IV Results and analysis
With the previously described setup, communication can be implemented. For simplicity, pre-determined messages can be encoded into the quantum states. In a secure implementation, an encoding scheme and random basis selection would be employed. The most straightforward message consists of a repeating string of alternating 0 s and 1 s throughout the communication. If the objective is solely to measure the QBER for each basis, the message can be encoded solely on the Z basis and then repeated on the X basis. To use the complete receiver, the basis must also be chosen.
@@ IV.1 Simplified receiver
The message is transmitted using the quantum states and detected using the HWP set to 0º for the Z basis and 22.5º for the X basis. Since the states are orthogonal, only one detector should register a click for each state sent. For the Z basis measurement, if the detector connected to the horizontal output of the PBS registers a click, the state sent was $|\mathrm{H}\rangle$, while a click in the other detector indicates the state was $|\mathrm{V}\rangle$. For educational purposes, it may be advantageous to define a higher average number of photons to ensure consistent detector clicks, simplifying the subsequent analysis.
@@
Assuming that bit 0 is transmitted by the horizontal polarization and bit 1 is transmitted by the vertical state, by analyzing which detector clicked more, a guess of the bit sent can be performed. If there are more events on channel 1, then the state sent was $|\mathrm{H}\rangle$, and therefore, the bit sent by that pulse was 0. The same logic for channel 2. By doing this to all pulses, a recovered message can be found.
By comparing this message to the original, counting the number of different bits, and dividing by the size of the message, the QBER can be found. We measure the QBER to be 2.5% with an uncertainty of 0.4% on the Z basis, while on the X basis, it is 2.11% with an uncertainty of 0.14%. These error rates are acceptable but could be reduced by addressing the impact of dark counts.
@@ IV.2 Complete receiver
If the complete receiver is used, the message sent has to have bits encoded in both bases. The procedure is similar to the simplified receiver, we obtain the data from the time tagger and have to recover the message. In this receiver, we have to recover which basis was chosen and the message sent. To recover the basis choice, we look at the number of events delayed and not delayed. If there are more delayed events, the basis chosen is X and if there are more photons without delay, the basis chosen is Z.
After, we can recover the message as done previously, looking into which detector clicked more. The basis choice can then be compared to Alice’s choice and only the events where there is a basis match are kept. Finally, the QBER is calculated in the same way, by comparing both messages.
This is a simplified way to estimate the QBER. Here the whole message is shared between Alice and Bob.
@@ IV.3 Simulating Eve
To simulate the presence of Eve one can create a very simple experiment. If we periodically block the free space channel during the communication phase, we can simulate an attack where Eve removes photons from the channel. This will lead to an increase in the QBER.
Another possible approach is to use a HWP to rotate the state polarization. If we use an HWP at 22.5º, we can simulate an Eve that always chooses the wrong measurement basis, increasing the QBER to around 50%. This gives intuition on an intercept and resend attempt where the basis choice is wrong all the time.
These experiments show that Eve’s presence in the channel leads to an increase in the QBER that can be detected aborting the communication and discarding the key generated as it was not secure.
@@ Appendix A Equipment price
Here we compile the prices and quantities of the essential components to implement the emitter and complete receiver. It does not include all the equipment necessary, ie, fibers, mechanical supports, and more.

### v2

@@ I Introduction
While educational articles exist on QKD implementation and single photon experiments in laboratories educational ; single ; double-slit ; letter , here we focus on the use of weak coherent pulses which offer a more straightforward and cost-effective setup, making quantum communication protocols accessible to undergraduate students. This work aims to bridge the gap between theory and practice, providing a valuable educational resource for understanding and implementing secure quantum communication.
@@ II.2 One Time Pad and BB84
Bob chooses one of the bases and obtains a bit from the result. If Alice and Bob chose the same basis, they will share the same bit value. For example, if Alice sends bit 0 in the Z basis and Bob performs a measurement on that basis, the probability of Bob measuring bit 0 is given by
@@
This means that Bob cannot recover the correct bit string when measuring on a different basis every round.
However, this measurement error, arising from a choice of different bases, can help us identify the presence of an eavesdropper (typically named Eve). If Eve intercepts and resends the state she measured from Alice, she will introduce errors.
As Eve does not know which basis Alice and Bob chose, she will guess a measurement basis. If she chooses a basis different from Alice and Bob, she will prepare a different state than the original. Thus, Eve will introduce errors.
@@
The error correction step is necessary to ensure that Alice and Bob share the same string of bits. This is accomplished using protocols like Cascade or LDPCs Error_corection , which require an estimation of the quantum bit error rate (QBER) and additional information to be shared through the classical channel (e.g., the parity of blocks of the key string). This estimation of the QBER is done by sharing part of the key between Alice and Bob. Finally, both parties run a privacy amplification protocol Renner to ensure that even with the extra information Eve might have intercepted, she cannot reconstruct the secret key. This procedure reduces the size of the key.
The performance of the communication protocol can then be evaluated by some key parameters like the QBER, the secret key rate (SKR), and the transmission distance (given by how much loss the system can tolerate before the SKR becomes negligible).
The QBER can be defined as the ratio of the number of erroneous bits to the total number of bits transmitted over a quantum channel. A lower QBER generally signifies a more secure and reliable QKD system, as it implies fewer errors in the transmitted key bits, increasing the potential for eavesdropping detection and the efficiency of the key generation process.
The SKR is the rate at which secure keys are generated and shared between two parties. It reflects the amount of usable key material produced per second after post-processing, ensuring secure communication. Higher secret key rates indicate more efficient and practical QKD systems.
The transmission distance is the distance at which the communication was performed and is related to the loss as for both free-space and fiber-based setups, the loss increases with the distance. The higher the loss an experimental setup can tolerate, the longer the transmission distance will be.
@@ II.3 Simplified BB84
[…]
For educational purposes focused on understanding the experimental implementation of secure communication through Quantum Key Distribution (QKD), a detailed description of error correction and privacy amplification was omitted. More information and the the security analysis can be found in .QBER
@@ III Experimental implementation
This setup is divided into two parts as seen in the appendix A, two optical breadboards of 60x60 cm (Thorlabs, B6060A), the emitter, and the receiver. These are the breadboards available in the laboratory but smaller breadboards could be used instead. The emitter is capable of sending weak coherent pulsesCoherent in three different polarization states ($|\mathrm{H}\rangle,\ |\mathrm{V}\rangle,\ |\mathrm{D}\rangle$). The receiver in turn is capable of measuring on two different bases, $Z$ and $X$.
@@ III.1 Emitter
The purple pulses are utilized to modulate the VCSEL to switch the laser on and off. The synchronization shown in the figure demonstrates the simultaneous modulation of these two types of pulses: blue for polarization control and purple for laser activation. This synchronization enables independent polarization control for each light pulse. In this case, the pulse duration was 10 $\mu s$ and the repetition rate was 50 kHz.
@@
Alternatively, it is also possible to implement a less costly approach, removing the EOM from the setup as seen in Fig. 3. Instead of one laser, three would be used. Only one laser would be turned on at any time and the path taken by the light would create the respective state. To create the $|\mathrm{H}\rangle$ and $\ |\mathrm{V}\rangle$ states, only a PBS would be required. To create the $|\mathrm{D}\rangle$ state, a PBS followed by a HWP can be used. This way, the state choice would be done by choosing which laser to turn on instead of by active modulation of the light. The different paths would then be recombined using a 50/50 beam splitter (BS)(Thorlabs, BS011) and the rest of the emitter would work as in the active approach.
This method brings security risks as the lasers can have manufacturing differences (spatial mode, frequency…) and an eavesdropper might take advantage of those but for this work, it can be a solution.
@@
Given that both the emitter and receiver are situated within the same room, the free-space channel is defined as the distance between the telescopes at each end. While a well-collimated beam and a short free-space channel may render the use of telescopes optional, their inclusion enables easy scalability of the channel length by adjusting the distance between optical tables. The emitter can be seen in Fig. 8 in the appendix A.
@@ III.2 Receivers
The objective of the receiver is to measure the incoming states on one of two basis. By calculating the QBER, we can detect the presence of potential eavesdroppers (Eve) and generate secret keys. For educational purposes focusing on the concept of key generation rather than actual key generation, this process can be simplified.
To accomplish this, as seen in Fig. 1, the light is captured by a telescope and directed through a HWP, followed by a PBS. The angle in the HWP will define if the measure is on the Z basis or the X basis. Each output of the PBS is then coupled to a fiber and a single-photon detector (SPD) (Excelitas, SPCM-AQRH-10). This detector was pre-aligned for multi-mode fibers as it simplifies the alignment but can be requested for single-mode pre-alignment.
The arrival time of the photons is recorded using a time tagger (Swabian Instruments, Time Tagger 20), which offers precision beyond the requirement; a more affordable option such as an electronic circuit or a field programmable gate array microcontroller can be used as described in. Coincidence ; Coincidence2 Additionally, the trigger pulse modulating the laser should be connected to the time tagger for post-processing of the measured data.
@@ III.3 Parameter definition
In a secure implementation of the protocol, the average number of photons emitted from the emitter should be low. However, for educational purposes, we may increase this value to simplify the setup, as even with large values of loss, the number of photons arriving at the detector can replicate what would be observed in a secure QKD experiment (single photon regime). This may leave the communication open to attacks.
This means that the use of a non-optimal fiber (or other sources of loss as misalignment) would still allow for the experimental implementation as it would only require an increase in the number of photons sent, to compensate for the extra loss.
@@
From Eq.(4), the necessary filter loss in dB can be determined. The parameters used in this work were: $\Delta T=10\mu s$, $P=92\mu W$, $\lambda=850$ nm, $\eta_{s}=12.2$ dB (including all the channel and receiver losses), and $\eta_{f}=80$ dB. These values result in a predicted $|\alpha|_{Alice}^{2}=39.3$ (weak coherent state emitted by Alice) and $|\alpha|_{Bob}^{2}=2.35$ (weak coherent state arriving at the detectors), while the measured value was $2.30\pm 0.05$. The measured probabilities of different photon counts were as follows: ”0” with a probability of 0.11, ”1” with 0.23, ”2” with 0.25, and ”3” with 0.19.
@@
To achieve this with only two detectors, measurements on the X basis are deliberately delayed. Despite the lack of refractive index information for the fiber used, typical values generally fall within the range of [1.45, 1.48]. Utilizing an available 1590-meter multi-mode optical fiber, although not ideal, yields an expected delay of approximately [7.685, 7.844] $\mu$s (around 5 ns per meter of fiber). As the number of photons can be adapted in this demonstration, the loss in the fiber is not a concern.
To verify this delay and assess time-multiplexed detection, a low-photon-count pulse with diagonal polarization is emitted from the emitter. For experimental validation, the BS in the setup (Fig 4) is replaced with a PBS to ensure minimal overlap between delayed and not delayed beams. The horizontal component of the polarized beam traverses a second PBS, reaching only detector 1 after passing through a filter to simulate losses in the alternate path, due to the fiber used. Conversely, the vertical component is directed into the optical fiber for temporal delay. Subsequently, a HWP rotates the polarization at the fiber output to horizontal, guaranteeing that photons exclusively reach detector 2 via a second PBS.
Consequently, clicks registered in detector 1 originate from the non-delayed signal, while those in detector 2 result from the delayed signal.
@@
The first normal distribution matches the expected outcome. As the pulse width is $10.0$ $\mu$s and the events follow a poisson distribution, there is a higher likelihood of registering an event at the center of the pulse width and zero probability of observing outside the pulse width, creating the observed normal distribution. The second distribution shows a broader width, likely due to dispersion in the fiber.
The observed overlap between the distributions can be attributed to selecting a delay shorter than the pulse width and to the broadening in the fiber, as now the normal distribution will have an event probability outside the original pulse width range. Despite this overlap, it remains minimal, allowing for clear differentiation between delayed and non-delayed signals, thereby facilitating the measurement of states in different bases.
@@ IV Results and analysis
With the previously described setup, communication can be implemented. For simplicity, pre-determined messages can be encoded into the quantum states. In a secure implementation, an encoding scheme and random basis selection and message generation would be employed. The most straightforward message consists of a repeating string of alternating 0 s and 1 s throughout the communication (”0101010101…”). For the evaluation of the setup, the QBER on each basis must be measured. To simplify this, the message can be encoded solely on the Z basis and then, the experiment is repeated on the X basis. For a more complete demonstration, the complete receiver can be used and the initial message has to be encoded on both bases, basis selection is necessary.
@@ IV.1 Simplified receiver
The message is transmitted using the quantum states and detected using the HWP set to 0º for the Z basis and 22.5º for the X basis. Since the states are orthogonal, only one detector should register a click for each state sent.
The single photon detectors create an electric pulse when an event is registered. This pulse will generate an event in the connected channel of the time tagger. This way, all the information is registered in the output file of the time tagger which includes the timestamp, the channel identification, and an error check parameter to access the correct use of the time tagger (missed events).
For the Z basis measurement, if the detector connected to the horizontal output of the PBS registers a click, the state sent was $|\mathrm{H}\rangle$, while a click in the other detector indicates the state was $|\mathrm{V}\rangle$. For educational purposes, it may be advantageous to define a higher average number of photons to ensure consistent detector clicks, simplifying the subsequent analysis. For the X basis, the same logic applies.
@@
Assuming that bit 0 is transmitted by the horizontal polarization and bit 1 is transmitted by the vertical state, by analyzing which detector clicked more, a guess of the bit sent can be performed. If there are more events on channel 1, then the state sent was $|\mathrm{H}\rangle$, and therefore, the bit sent by that pulse was 0. The same logic applies to channel 2. By doing this to all pulses, a recovered message can be found.
Using this logic, from Fig. 6, the string ”010” is recovered, which matches the bits sent for this fraction of the message. By collecting data for 10 seconds, a message of 500 kbit is sent by Alice and recovered by Bob. By comparing this message to the original, counting the number of different bits, and dividing by the size of the message, the QBER can be found. We measure the QBER to be 2.5% with an uncertainty of 0.4% on the Z basis, while on the X basis, it is 2.11% with an uncertainty of 0.14%. These error rates are acceptable but could be reduced by addressing the impact of dark counts.
@@
While this implementation did not focus on optimizing the SKR, it is still an important figure of merit. As we did not implement an error correction code and privacy amplification protocol we estimated the secret key size using the upper bound provided in.SimpleBB84
As we used a weak coherent state with a high number of photons, the corresponding secret key rate is 0. The communication is open to attacks as Eve can capture a few photons each round and recover the entire message. For a secure implementation, the number of photons sent by Alice has to be reduced, and by using a python package created to estimate the SKR of the simplified BB84 for space-based communication, provided in [^1] The code is available on the Github page https://github.com/QuLab-IT/QuantSatSimulator.git., we arrived at the optimal secret key rate of 75 bps for an average photon number of 0.87 and parameters reported in the appendix B. This value can be easily improved by increasing the source rate or the communication time window.
[…]
Recent works have shown a SKR of 115.8 Mbps over 10-km standard fiber for a source of 2.5 GHz, proving that high-rate quantum key distribution is possible high_rate . There are also commercially available products like the Clavis XGR QKD platform from ID Quantique operating with a 1 Gbit source and generating secret keys at around 12 dB of loss at a rate of 400 kbps. [^2] Clavis XGR QKD https://www.idquantique.com/quantum-safe-security/products/clavis-xgr-qkd-platform/.
@@ IV.2 Complete receiver
If the complete receiver is used, the message sent has to have bits encoded in both bases, closer to what is used in a secure QKD setup. The procedure is similar to the simplified receiver, we obtain the data from the time tagger and must recover the message. In this receiver, we have to recover which basis was chosen and the message sent. To recover the basis choice, we look at the number of events delayed and not delayed. This means that the timestamp data shown in Fig. 6 has to be used. If there are more delayed events, events 10 $\mu s$ after an event in channel 4, the basis chosen is X, and if there are more photons without delay, events between the timestamp of channel 4 and 10 $\mu s$ after, the basis chosen is Z.
After, we can recover the message as done previously, looking into which detector clicked more. The basis choice can then be compared to Alice’s choice and only the events with a basis match are kept. Finally, the QBER is calculated in the same way, by comparing both messages.
This is a simplified way to estimate the QBER where the whole message is shared between Alice and Bob, not just a fraction. This is followed again by the post-processing to generate the secret key.
@@ IV.3 Simulating Eve
To simulate the presence of Eve one can create a very simple experiment. If we periodically block the free space channel during the communication phase, we can simulate an attack where Eve removes photons from the channel. This will stop all detection, leading to an increase in the QBER. For example, if for no detection we assume bit 0 was sent, the message recovered while Eve is blocking the channel would be an infinite string of 0s.
Another possible approach is to use a HWP to rotate the state polarization. If we use an HWP at 22.5º, we can simulate an Eve that always chooses the wrong measurement basis, increasing the QBER to around 50% as the message recovered would be random (the measurement is made on X basis while the state was encoded on the Z basis for example). This gives intuition on an intercept and resend attack where the basis choice is always wrong. The change in the experimental setup can be seen in Fig. 7.
Figure 7: Design of Eve simulation experiments. a) The free space channel is blocked. b) A HWP is used to rotate the basis.
These simple experiments show that Eve’s presence in the channel leads to an increase in the QBER that can be detected aborting the communication and discarding the key generated as it was not secure.
V Conclusion
In summary, this work has demonstrated a practical, compact, and cost-effective QKD setup suitable for undergraduate tutorial demonstrations. By using weak coherent pulses, the simplified three-state BB84 protocol, and time multiplexing, the complexity of the setup can be reduced and QKD demonstrations can be easily implemented, making the setup more accessible and manageable within educational environments.
@@ Appendix A Equipment Price
Here we compile the prices and quantities of the essential components to implement the emitter and complete receiver. It does not include all the equipment necessary, i.e., fibers, mechanical supports, and more.
@@
Appendix B Parameters for SKR Optimization
The parameters used for the optimization of the SKR for this setup is shown in Table 3.
Parameter Value
Signal intensity 0.79
Probability of sending signal 0.75
Decoy intensity 0.02
Probability Alice sends an $Z$ basis signal 0.5
Probability Bob measures an $Z$ basis signal 0.5
Loss 12 dB
Communication time 10 s
Correctness parameter ($\epsilon_{C}$) $10^{-15}$
Secrecy parameter ($\epsilon_{S}$) $10^{-9}$
Intrinsic Quantum Bit Error Rate $10^{-3}$
Extraneous count probability $10^{-3}$
After pulse probability $10^{-3}$
Source repetition rate 100 kHz
Table 3: Parameters used for SKR optimization